Solve the quetion according to "Volume and Surface Area" formula
A rectangular tank measures 4m long,2m wide and 4.8m high.Intially it is half filled with water .Find the depth of water in the tank after 4000 liters more of water added to it .

Answer by mananth(16946)

A rectangular tank measures 4m long,2m wide and 4.8m high.Intially it is half filled with water .Find the depth of water in the tank after 4000 liters more of water added to it .
..
Volume of full tank = 4*2*4.8 =38.4 cu.m
1/2 volume = 38.4 /2 = 19.2 cu.m
19.2 cu. m = 19.2 *1000 liters
= 19200 liters
..
19200 liters = 2.4 m height ( half tank)
4000 liters = ?
= 2.4*4000/19200
=0.5 meters height for 4000 l
Total depth will be 2.4 + 0.5= 2.9 m